jagr2713 (jagr2713):

f(7)= -x2+3x-6 i got 64 but its says its wrong

OpenStudy (freckles):

\[f(x)=-x^2+3x-6 \\ \text{ and you are asked to evaluate } f(7) ?\]

OpenStudy (freckles):

by the way -x^2 means -1*x*x

OpenStudy (freckles):

-x^2 is not the same as (-x)^2

OpenStudy (freckles):

which looks like what you did

jagr2713 (jagr2713):

oh it does 0_0 i put -7^2+3(7)-6

OpenStudy (freckles):

which is -49+21-6

jagr2713 (jagr2713):

wait for mines i got 49+21-6 = 64

OpenStudy (freckles):

as said you did (-x)^2

OpenStudy (freckles):

but the question says -x^2

OpenStudy (freckles):

-x^2 is -1*x*x not (-x)*(-x)

OpenStudy (freckles):

\[-x^2=-1 \cdot x^2 \text{ or } -1 \cdot x \cdot x \\ \text{ where as } (-x)^2 =(-x) \cdot (-x)=x \cdot x =x^2 \text{ which is what you did } \\ \text{ and } -x^2 \text{ does not equal } (-x)^2 \]

jagr2713 (jagr2713):

so can u give it to me when you first plug it in

jagr2713 (jagr2713):

so its -7^2 ?

OpenStudy (freckles):

yes not (-7)^2

OpenStudy (freckles):

the exponent is on the x not the - and the x

OpenStudy (freckles):

\[g(x)=-x^2 \\ g(7)=-7^2=-(7^2)=-(7 \cdot 7)=-(49)=-49\]

jagr2713 (jagr2713):

so thats -49?

OpenStudy (freckles):

if you had \[h(x)=(-x)^2 \\ \text{ then } h(7)=(-7)^2=(-7 \cdot -7)=49\]

jagr2713 (jagr2713):

OHOHOHOHHOHO i seeeeeee

jagr2713 (jagr2713):

so if its for example f(8) we put -8^2 so 8*8=64 and the negative sign is after so -64

OpenStudy (freckles):

yes -8^2 means -(8*8)=-(64)=-64

OpenStudy (freckles):

where as (-8)^2=(-8)*(-8)=64

jagr2713 (jagr2713):

so the answer is *34?

OpenStudy (freckles):

do you see difference between -x^2 and (-x)^2 ?

jagr2713 (jagr2713):

-34?

OpenStudy (freckles):

\[-7^2+3(7)-6 \\ -49+21-6 \\ -(49-21)-6 \\ -(28)-6 \\ -28-6 \\ -34\] yes

jagr2713 (jagr2713):

OMG your the best thanks!!!!!

OpenStudy (freckles):

np